### v2

# Certified Resilient Collision-free Control Framework for Heterogeneous Multi-agent Systems Against Exponentially Unbounded FDI Attacks

Abstract
False data injection attacks pose a significant threat to autonomous multi-agent systems (MASs). Existing attack-resilient control strategies generally have strict assumptions on the attack signals and overlook safety constraints, such as collision avoidance. In practical applications, leader agents equipped with advanced sensors or weaponry span a safe region to guide heterogeneous follower agents, ensuring coordinated operations while addressing collision avoidance to prevent financial losses and mission failures. This letter addresses these gaps by introducing and solving the safety-aware and attack-resilient (SAAR) control problem under exponentially unbounded false data injection (EU-FDI) attacks. Specifically, a novel attack-resilient observer layer (OL) is first designed to defend against EU-FDI attacks on the OL. Then, an attack-resilient compensational signal is designed to mitigate the adverse effects caused by the EU-FDI attack on control input layer (CIL). Finally, a SAAR controller is designed by solving a quadratic programming (QP) problem integrating control barrier function (CBF) certified collision-free safety constraints. Rigorous Lyapunov-based stability analysis certifies the SAAR controller’s effectiveness in ensuring both safety and resilience. This study also pioneers a three-dimensional (3D) simulation of the SAAR containment control problem for heterogeneous MASs, demonstrating its applicability in realistic multi-agent scenarios.
I Introduction
Containment control in autonomous MASs has become essential in real-world applications, particularly involving unmanned aerial vehicles (UAVs) [1] and unmanned ground vehicles (UGVs) [2]. This control strategy enables leader agents, often equipped with advanced sensors or weaponry, to guide follower agents within a designated safe region, ensuring coordinated movement and operational safety. Such scenarios are common in tasks such as surveillance, reconnaissance, and military operations, where maintaining cohesion and mitigating risks are critical. The heterogeneity of agents, characterized by differing models, structures, or capabilities, adds complexity to the problem, necessitating advanced control algorithms. Moreover, collision avoidance plays a pivotal role in containment control, as collisions can lead to substantial financial losses and mission failures, highlighting the importance of integrating safety measures into control strategies.
In the MAS framework, FDI attacks pose significant threats to autonomous MASs and impact critical infrastructures and mission-essential cyber-physical systems, such as power systems, transportation networks, and combat-zone multi-robot systems [3]. To mitigate the impact of FDI attacks, two primary approaches have been proposed. The first one involves detection and identification of compromised agents, followed by the attack-isolation [4, 5]. However, this method relies on strict limitations, such as the number of compromised agents, which are often impractical. To address these limitations, a second approach has been developed, focusing on the design of attack-resilient control protocols [6, 7, 8, 9, 10]. Rather than detecting and removing compromised agents, this strategy aims to minimize the adverse effects of attacks through resilient control mechanisms. However, the aforementioned research neglects safety constraints, such as collision avoidance, leading to potential collision risks. On the other hand, existing safe control methods incorporating safety constraints often fail to guarantee resilience for MASs against FDI attacks [11, 12, 13, 14, 15]. Additionally, most studies addressing FDI resilience for heterogeneous MASs assume the OL remains intact and either disregard attacks or handle only bounded attack signals or the ones with bounded first-time derivatives, which are impractical assumptions [16, 7]. However, malicious adversaries could inject any time-varying signals, such as EU-FDI attacks, to compromise the system.
In real-world scenarios, ensuring both safety and resilience against EU-FDI attacks is crucial for the reliable operation of safety-critical MASs. To address this challenge, this letter formulates and solves the SAAR control problem. The key difficulty lies in designing a controller that effectively mitigates EU-FDI attacks while consistently enforcing collision avoidance constraints. The main contributions are as follows.
$\bullet$ A novel attack-resilient OL is first designed to defend against EU-FDI attacks on OL. Then, an attack-resilient compensation signal is developed to mitigate the adverse effects of EU-FDI attacks on CIL. Finally, the SAAR controller is designed by solving a QP problem incorporating CBF-certified collision-free safety constraints. To the best of the authors’ knowledge, this letter is the first to guarantee both resilience and safety for heterogeneous MASs, in the presence of EU-FDI attacks on both OL and CIL.
$\bullet$ A rigorous Lyapunov-based stability analysis is conducted to certify the effectiveness of the proposed SAAR controller in ensuring both safety and resilience. Its efficacy is further demonstrated through a 3D simulation of the containment control problem for heterogeneous MASs.
II Preliminaries and Problem Formulation
$\|x\|$ denotes the Euclidean norm of a vector $x\in\mathbb{R}^{n}$.
$L_{f}V(x)=\frac{\partial V}{\partial x}f(x)$ is the Lie derivative of a function $V:\mathbb{R}^{n}\rightarrow\mathbb{R}$ along a vector field $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. ${I_{N}}\in{\mathbb{R}^{N\times N}}$ is the identity matrix. ${\mathbf{1}}_{N},{\mathbf{0}}_{N}\in{\mathbb{R}^{N}}$ are the column vectors with all elements of zero and one, respectively. The Kronecker product is represented by $\otimes$. The operator $\operatorname{diag}(\cdot)$ forms a block diagonal matrix from its argument. The notations ${\sigma_{\min}}(\cdot)$ and ${\sigma_{\max}}(\cdot)$ represent the minimum singular value and the maximum singular value respectively. Consider a system of $N+M$ agents on a time-invariant digraph $\mathscr{G}$, consisting of $N$ followers and $M$ leaders, denoted by =$\mathscr{F}=\{v_{1},v_{2},\ldots,v_{N}\}$ and =$\mathscr{L}=\{v_{N+1},v_{N+2},\ldots,v_{N+M}\}$, respectively. $\mathcal{N}_{i}$ represents the set of neighboring followers of follower $i$. The interactions among followers are described by the subgraph $\mathscr{G}_{f}=(\mathcal{V},\mathcal{E},\mathcal{A})$. More details on the digraph can be found in [6]. Denote $\Phi_{r}=\frac{1}{M}{\mathcal{L}}+{\mathcal{G}}_{r}$. The states of the leaders are represented by the set $X_{\mathscr{L}}=\{x_{N+1},x_{N+2},\ldots,x_{N+M}\}$.
We consider a group of $N$ heterogeneous followers described by the following dynamics:
$\dot{x}_{i}(t)=A_{i}x_{i}(t)+B_{i}u_{i}(t),\quad i\in\mathscr{F},$ (1)
where $x_{i}(t)\in\mathbb{R}^{n}$ is the state, $u_{i}(t)\in\mathbb{R}^{m_{i}}$ is the control input. The system matrices $A_{i}$ and $B_{i}$ may vary across different agents, making the system heterogeneous. The $M$ leaders with the following dynamics can be viewed as command generators that generate the desired trajectories:
$\dot{x}_{r}(t)=Sx_{r}(t),\quad r\in\mathscr{L},$ (2)
where $x_{r}(t)\in\mathbb{R}^{n}$ is the state of the $r$th leader.
Definition 1 ([17]).
The convex hull $\text{Co}(X_{\mathscr{L}})$ is the minimal convex set containing all points in $X_{\mathscr{L}}$, defined as:
$\text{Co}(X_{\mathscr{L}})=\left\{\sum\nolimits_{r\in\mathscr{L}}a_{r}x_{r}\;\middle|\;a_{r}\geqslant 0,\sum\nolimits_{r\in\mathscr{L}}a_{r}=1\right\}.$
where $\sum_{r\in\mathscr{L}}a_{r}x_{r}$ represents certain convex combination of all points in $X_{\mathscr{L}}$.
Definition 2 ([18]).
The signal $x(t)\in{\mathbb{R}^{n}}$ is said to be UUB with the ultimate bound $b$, if there exist positive constants $b$ and $c$, independent of ${t_{0}}\geqslant 0$, and for every $a\in\left({0,c}\right)$, $\exists\,T=T\left({a,b}\right)\geqslant 0$, independent of $t_{0}$, such that
$\left\|{x\left({{t_{0}}}\right)}\right\|\leqslant a\;\;\Rightarrow\;\;\left\|{x\left(t\right)}\right\|\leqslant b,\forall t\geqslant{t_{0}}+T$ (3)
Assumption 1.
Each follower in the digraph ${\mathscr{G}}$, has a directed path from at least one leader.
Assumption 2.
$S$ has eigenvalues with non-positive real parts, and non-repeated eigenvalues on the imaginary axis.
Assumption 3.
The pair $(A_{i},B_{i})$ is controllable for each follower.
Given Assumption 3, the following linear matrix equation has a solution $\Pi_{i}$ for each follower:
$S=A_{i}+B_{i}\Pi_{i}.$ (4)
Assumption 4.
The FDI attack signals on the CIL and OL, $\gamma^{a}_{i}(t)$ and $\gamma^{ol}_{i}(t)$, are exponentially unbounded. That is, their norms grow at most exponentially with time. For the purposes of stability analysis, it is reasonable to assume that there exist positive constants $\kappa^{a}_{i}$ and $\kappa^{ol}_{i}$, such that $\|\gamma^{a}_{i}(t)\|\leqslant\exp(\kappa^{a}_{i}t)$ and $\|\gamma^{ol}_{i}(t)\|\leqslant\exp(\kappa^{ol}_{i}t)$, where $\kappa^{a}_{i}$ and $\kappa^{ol}_{i}$ could be unknown.
Remark 1.
Assumption 4 covers a broad range of FDI attack signals, including those that grow exponentially over time. Specifically, $\exp(\kappa^{a}_{i}t)$ and $\exp(\kappa^{ol}_{i}t)$ represent the worst-case scenarios the controller can handle. As long as the attack signals remain below these envelops, the controller can effectively mitigate them. In practice, adversaries can inject any time-varying signal into MAS through platforms such as software, CPUs, or DSPs. However, most of the existing research focuses on disturbances, noise, or bounded attack signals, or assumes that the first time derivatives of the attacks are bounded [7].
Let $\bar{x}_{r}=\mathbf{1}_{N}\otimes x_{r}$. Define the global containment error as
$\begin{gathered}e_{c}=x-\left(\sum\nolimits_{\nu\in\mathscr{L}}\left(\Phi_{\nu}\otimes I_{n}\right)\right)^{-1}\sum\nolimits_{r\in\mathscr{L}}\left(\Phi_{r}\otimes I_{n}\right)\bar{x}_{r},\end{gathered}$ (5)
where $x={[{x_{1}^{\top},...,x_{N}^{\top}}]^{\top}}$.
Lemma 1 ([19]).
Given Assumption 1, the containment control objective is achieved if $\lim_{t\to\infty}e_{c}(t)=\mathbf{0}$. That is, each follower converges to the convex hull spanned by the leaders.
Definition 3.
A set $S\subset\mathbb{R}^{n}$ is forward invariant with respect to a control system if, for all initial conditions $x(0)\in S$, the system’s trajectory satisfies $x(t)\in S$ for all $t\geqslant 0$.
The forward invariance of the safe set is guaranteed using CBFs, as described as follows.
Definition 4 ([20]).
Let $\alpha$ be a monotonically increasing, locally Lipschitz class $\mathcal{K}$ function with $\alpha(0)=0$, given a set $S\subset\mathbb{R}^{n}$ as defined in Definition 3, a function $h(x)$ is CBF, if the following condition is satisfied:
$\inf_{u\in\mathcal{U}}\left[L_{f_{i}}h(x)+L_{g_{i}}h(x)u\right]\leqslant-\alpha(h(x)),$ (6)
where $\mathcal{U}$ is the set of admissible control inputs.
In (6), $L_{f}h(x)$ and $L_{g}h(x)$ denote the Lie derivatives of the function $h(x)$ along the vector fields $f$ and $g$, respectively. For a general system of the form: $\dot{x}=f(x)+g(x)u,$ the terms $L_{f}h(x)=\frac{\partial h}{\partial x}f(x)$ and $L_{g}h(x)=\frac{\partial h}{\partial x}g(x)$ involve the gradients of $h(x)$ with respect to the state $x$. if control input $u$ is designed such that (6) is satisfied, the forward invariant set $S$ is obtained [21].
Now, consider our specific case where $h(x)$ applied locally corresponds to $h_{S_{ij}}(x_{i},x_{j})$. We apply aforementioned general concepts to our specific case, where the safe set for collision avoidance is defined as:
$S_{S_{ij}}\triangleq\{x_{i}\in\mathbb{R}^{n}\mid h_{S_{ij}}(x_{i},x_{j})\leqslant 0\},$ (7)
with the specific CBF:
$h_{S_{ij}}(x_{i},x_{j})\triangleq d_{s}^{2}-\|x_{i}-x_{j}\|^{2},$ (8)
where $d_{s}$ is the minimum allowable safe distance. The system dynamics are given by: $\dot{x}_{i}(t)=A_{i}x_{i}(t)+B_{i}u_{i}(t),\,\forall i\in\mathscr{F}.$ In this case, the vector field $f_{i}(x)$, applied to linear time invariant system locally, corresponds to $A_{i}x_{i}$, and $g_{i}(x)$ corresponds to $B_{i}$. To avoid confusion, we denote the Lie derivatives as $L_{f_{i}}h_{S_{ij}}$ and $L_{g_{i}}h_{S_{ij}}$, where $f_{i}$ and $g_{i}$ are defined as above.
The following definition introduces the SAAR control problem.
Definition 5 (SAAR Control Problem).
For the heterogeneous MAS described in (1)-(2) under EU-FDI attacks, the SAAR control problem is to design control input $u_{i}\in\mathcal{U}_{i}$, such that 1) $e_{c}$ is UUB, i.e., the UUB containment control objective is achieved. Specifically, the state of each follower converges to a small neighborhood around or within the dynamic convex hull spanned by the states of the leaders; 2) the state $x_{i}(t)$ remains within the safe set $S_{S_{ij}}$ for all $t\geqslant 0$ in the presence of EU-FDI attacks on both OL and CIL.
III Main Result of SAAR Controller Design
In this section, we propose a fully distributed, safe, and resilient containment control framework to address the SAAR control problem. Since only the neighboring followers of the leaders have access to the leaders’ states, to achieve the containment control objective for each follower, a OL dynamics is needed to estimate the convex combinations of the leaders’ states. Here, we introduce the following fully distributed and attack-resilient OL dynamics
$\dot{\zeta}_{i}=S\zeta_{i}+\exp(\vartheta_{i})\xi_{i}+\gamma^{ol}_{i},$ (9)
$\dot{\vartheta}_{i}=q_{i}\xi_{i}^{\top}\xi_{i},$ (10)
where $\zeta_{i}$ is the local state on the OL and $\gamma^{ol}_{i}$ represents the attack signal on the OL, $\vartheta_{i}$ is adaptively tuned by (10) with constant $q_{i}>0$, and $\xi_{i}$ represents the gathered neighborhood relative information on the OL given by
$\xi_{i}=\sum_{j\in\mathscr{F}}a_{ij}(\zeta_{j}-\zeta_{i})+\sum_{r\in\mathscr{L}}g_{ir}(x_{r}-\zeta_{i}),i\in\mathscr{F}$ (11)
Based on this OL, we then introduce the following conventional control input design [19].
$\begin{gathered}u^{c}_{i}=K_{i}x_{i}+H_{i}\zeta_{i},\hfill\end{gathered}$ (12)
The matrices $K_{i}$ and $H_{i}$ are obtained as:
$K_{i}=-U_{i}^{-1}B_{i}^{\top}P_{i},$ (13)
$H_{i}=\Pi_{i}-K_{i},$ (14)
where the matrix $P_{i}$ is obtained by solving the following algebraic Riccati equation[7]:
$A_{i}^{\top}P_{i}+P_{i}A_{i}+Q_{i}-P_{i}B_{i}U_{i}^{-1}B_{i}^{\top}P_{i}=0.$ (15)
Define the follower-observer tracking error as
$\varepsilon_{i}=x_{i}-\zeta_{i}.$ (16)
We consider the additional EU-FDI attack signal in the CIL for each follower, $\gamma^{a}_{i}$. To address these attacks, we then design the following attack-resilient control input signal
$u^{r}_{i}$ $=u^{c}_{i}-\hat{\gamma}^{a}_{i},$ (17)
$\hat{\gamma}^{a}_{i}$ $=\frac{B_{i}^{\top}P_{i}\varepsilon_{i}}{\|\varepsilon_{i}^{\top}P_{i}B_{i}\|+\exp(-c_{i}t^{2})}\exp(\hat{\rho}_{i}),$ (18)
$\dot{\hat{\rho}}_{i}$ $=\alpha_{i}\|\varepsilon_{i}^{\top}P_{i}B_{i}\|,$ (19)
where the compensational signal $\hat{\gamma}^{a}_{i}$ mitigates the adverse effects caused by the attack $\gamma^{a}_{i}$, and $\hat{\rho}_{i}$ is adaptively tuned by (19) with constant $\alpha_{i}>0$. Now the corrupted attack-resilient controller becomes $\bar{u}_{i}=u^{r}_{i}+\gamma^{a}_{i}$.
Define the observer containment error vector as
$\begin{gathered}\Delta_{o}=\zeta-\left(\sum_{\nu\in\mathscr{L}}\left(\Phi_{\nu}\otimes I_{n}\right)\right)^{-1}\sum_{r\in\mathscr{L}}\left(\Phi_{r}\otimes I_{n}\right)\bar{x}_{r},\end{gathered}$ (20)
Lemma 2 ([19]).
Given Assumption 1, $\sum_{r\in\mathscr{L}}{\Phi_{r}}$ is non-singular and positive-definite.
Subsequently, we integrate the attack-resilience with collision-free safety as follows.
Lemma 3.
Given the set $S_{S_{ij}}$ and the CBF $h_{S_{ij}}(x_{i},x_{j})$, if the control input $u_{i}$ satisfies:
$\begin{gathered}\inf_{u_{i}\in\mathcal{U}_{i}}\left[L_{f_{i}}h_{S_{ij}}+L_{g_{i}}h_{S_{ij}}u_{i}\right]\leqslant-\alpha(h_{S_{ij}})\hfill\\ -2(x_{i}-x_{j})^{\top}A_{j}x_{j}-2(x_{i}-x_{j})^{\top}B_{j}u_{j},\end{gathered}$ (21)
then the system ensures that $\|x_{i}-x_{j}\|\geqslant d_{s}$ for all $t\geqslant 0$.
Proof: By definition, the specific CBF is given as: $h_{S_{ij}}(x_{i},x_{j})=d_{s}^{2}-\|x_{i}-x_{j}\|^{2}.$ Taking the time derivative of $h_{S_{ij}}(x_{i},x_{j})$ along the system dynamics, we have: $\dot{h}_{S_{ij}}(x_{i},x_{j})=-\frac{d}{dt}\|x_{i}-x_{j}\|^{2}.$ Expanding $\|x_{i}-x_{j}\|^{2}=(x_{i}-x_{j})^{\top}(x_{i}-x_{j})$, its time derivative is $\frac{d}{dt}\|x_{i}-x_{j}\|^{2}=2(x_{i}-x_{j})^{\top}(\dot{x}_{i}-\dot{x}_{j}).$ Substituting this into $\dot{h}_{S_{ij}}(x_{i},x_{j})$, we obtain: $\dot{h}_{S_{ij}}(x_{i},x_{j})=-2(x_{i}-x_{j})^{\top}(\dot{x}_{i}-\dot{x}_{j}).$ From the system dynamics, we know: $\dot{x}_{i}=A_{i}x_{i}+B_{i}u_{i},\quad\dot{x}_{j}=A_{j}x_{j}+B_{j}u_{j}.$ Substituting $\dot{x}_{i}$ and $\dot{x}_{j}$ into the expression for $\dot{h}_{S_{ij}}(x_{i},x_{j})$, we get: $\dot{h}_{S_{ij}}(x_{i},x_{j})=-2(x_{i}-x_{j})^{\top}\left[(A_{i}x_{i}+B_{i}u_{i})-(A_{j}x_{j}+B_{j}u_{j})\right].$ Using the Lie derivative notation, this can be written as: $\dot{h}_{S_{ij}}(x_{i},x_{j})=L_{f_{i}}h_{S_{ij}}+L_{g_{i}}h_{S_{ij}}u_{i}+2(x_{i}-x_{j})^{\top}A_{j}x_{j}+2(x_{i}-x_{j})^{\top}B_{j}u_{j},$ where, $L_{f_{i}}h_{S_{ij}}=-2(x_{i}-x_{j})^{\top}A_{i}x_{i},L_{g_{i}}h_{S_{ij}}=-2(x_{i}-x_{j})^{\top}B_{i}$. Based on [21], to ensure forward invariance of $S_{S_{ij}}$, we impose the condition:
$\dot{h}_{S_{ij}}(x_{i},x_{j})\leqslant-\alpha(h_{S_{ij}}).$
Substituting the expanded form of $\dot{h}_{S_{ij}}(x_{i},x_{j})$, we obtain: $L_{f_{i}}h_{S_{ij}}+L_{g_{i}}h_{S_{ij}}u_{i}\leqslant-\alpha(h_{S_{ij}})-2(x_{i}-x_{j})^{\top}A_{j}x_{j}-2(x_{i}-x_{j})^{\top}B_{j}u_{j}.$ This inequality ensures that the safe set $S_{S_{ij}}$ is forward invariant. Thus, if the control input $u_{i}$ satisfies the optimization criterion:
$\inf_{u_{i}\in\mathcal{U}_{i}}\left[L_{f_{i}}h_{S_{ij}}+L_{g_{i}}h_{S_{ij}}u_{i}\right]\leqslant-\alpha(h_{S_{ij}})-2(x_{i}-x_{j})^{\top}A_{j}x_{j}-2(x_{i}-x_{j})^{\top}B_{j}u_{j},$ then $\|x_{i}-x_{j}\|\geqslant d_{s}$ is guaranteed for all $t\geqslant 0.\hfill\blacksquare$
Figure 1: Pairs.
For control affine systems, CBFs lead to linear constraints on the control inputs $u_{i}$ that can be enforced online. To ensure safety constraints and prevent collisions, we design the SAAR control input $u_{i}$ by solving the following QP. This optimization minimizes the difference between the actual SAAR control and the corrupted attack-resilient control input $\bar{u}_{i}$, while enforcing the constraints:
$\min_{u_{i}}\|u_{i}-\bar{u}_{i}\|^{2},$ (22)
subject to the constraints:
$\begin{gathered}\inf_{u_{i}\in\mathcal{U}_{i}}\left[L_{f_{i}}h_{S_{ij}}+L_{g_{i}}h_{S_{ij}}u_{i}\right]\leqslant-\delta_{ij}h_{S_{ij}}\hfill\\ -2(x_{i}-x_{j})^{\top}A_{j}x_{j}-2(x_{i}-x_{j})^{\top}B_{j}u_{j},\end{gathered}$ (23)
where $1\leqslant i<j<N$, which ensures the pairwise distances are defined for all agent pairs $(i,j)$ and each pair is counted exactly once as shown in Fig 1. $\delta_{ij}$ is the parameters which influences the intensity of the safety constraints. We select $\alpha(h_{S_{ij}})=\delta_{ij}h_{S_{ij}}$[22]. Let $\Delta u_{i}=u_{i}-\bar{u}_{i}$, which is obtained by solving a convex QP with linear constraints over the feasible set $\mathcal{U}_{i}$ (which is assumed to be compact or otherwise bounded in practice). Therefore, $\Delta u_{i}$ is bounded. The efficient solution of the QP problem enables the algorithm to be implemented in real-time [23]. Collision avoidance among agents is enforced through a sequential backward-constrained optimization scheme. Starting from the highest-indexed agent pair, the solution for each agent $i$ is computed using constrained optimization based on results from previously solved pairs $(i,j)$, where $j>i$. For example, if the total number of followers $N=4$, the resulting pairs are $(1,2)$, $(1,3)$, $(1,4)$, $(2,3)$, $(2,4)$, and $(3,4)$, fully enumerating all distinct pairwise interactions. In the updating algorithm, inputs for agents are sequentially updated by first addressing pairs involving the highest-indexed agents. Specifically, the input of the highest-indexed agent remains unmodified, i.e., $u_{4}=\bar{u}_{4}$, the input for agent 3 is updated based on the $(3,4)$ pair. Then, using the updated information for agents 3 and 4, the input for agent 2 is updated through pairs $(2,3)$ and $(2,4)$. Finally, inputs for agent 1 are updated simultaneously using the pairs $(1,2)$, $(1,3)$, and $(1,4)$, incorporating previously updated inputs. An alternative approach from [21] with reduced neighborhoods on a disk graph provides a scalable solution with less computational cost.
The proposed SAAR control system is shown in Fig. 2.
Figure 2: Block diagram of the SAAR control system.
Theorem 1.
Given Assumptions 1- 4, considering the heterogeneous MAS composed of (1)-(2) in the presence of EU-FDI attacks on both CIL and OL, the SAAR control problem is solved by designing the fully-distributed safe and resilient controller consisting of (9)-(19) and (22)-(23).
Proof: The global form of (11) is
$\begin{gathered}\xi=-\sum_{\nu\in\mathscr{L}}(\Phi_{\nu}\otimes I_{N})\big{(}\zeta-\bar{x}_{r}\big{)}=-\sum_{\nu\in\mathscr{L}}(\Phi_{\nu}\otimes I_{N})\Delta_{o},\hfill\end{gathered}$ (24)
where $\xi={[{\xi_{1}^{\top},...,\xi_{N}^{\top}}]^{\top}}$. Based on Lemma 2, $\sum_{\nu\in\mathscr{L}}(\Phi_{\nu}\otimes I_{n})$ is nonsingular. To prove $\Delta_{o}$ is UUB is equivalent to proving that $\xi$ is UUB. The global form of $\dot{\zeta}_{i}$ in (9) is
$\begin{gathered}\dot{\zeta}=(I_{N}\otimes S)\zeta+\operatorname{diag}(\exp(\vartheta_{i}))\xi+\gamma^{ol},\hfill\end{gathered}$ (25)
where $\gamma^{ol}={[{{\gamma^{ol}_{1}}^{\top},...,{\gamma^{ol}_{N}}^{\top}}]^{\top}}$. The time derivative of $\xi$ in (24) is
$\begin{gathered}\dot{\xi}=-\sum\nolimits_{\nu\in\mathscr{L}}(\Phi_{\nu}\otimes I_{n})\big{(}(I_{N}\otimes S)\zeta+\hfill\\ \big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{n}}\big{)}\xi+\gamma^{ol}-(I_{N}\otimes S)\bar{x}_{r}\big{)}\hfill\\ =(I_{N}\otimes S)\xi-\sum_{r\in\mathscr{L}}(\Phi_{r}\otimes I_{n})\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{n}}\big{)}\xi\hfill\\ -\sum\nolimits_{r\in\mathscr{L}}(\Phi_{r}\otimes I_{n})\gamma^{ol}.\hfill\end{gathered}$ (26)
We consider the following Lyapunov function candidate
$V^{{}^{\prime}}=\tfrac{1}{2}\sum\nolimits_{i\in\mathscr{F}}\xi_{i}^{\top}\xi_{i}\exp(\vartheta_{i}).$ (27)
The time derivative of $V^{{}^{\prime}}$ along the trajectory of (26) is
$\begin{gathered}\dot{V}^{{}^{\prime}}=\sum\nolimits_{i\in\mathscr{F}}\big{(}\xi_{i}^{\top}{\dot{\xi}_{i}}\exp(\vartheta_{i})+\frac{1}{2}\xi^{T}_{i}\xi_{i}\exp(\vartheta_{i})\dot{\vartheta}_{i}\big{)}\hfill\\ =\xi^{\top}\operatorname{diag}\big{(}\exp(\vartheta_{i})\otimes I_{N}\big{)}\dot{\xi}+\frac{1}{2}\xi^{T}_{i}\big{(}\operatorname{diag}(\exp(\vartheta_{i})\dot{\vartheta}_{i})\otimes I_{N}\big{)}\xi\hfill\\ =\xi^{\top}\operatorname{diag}\big{(}\exp(\vartheta_{i})\otimes I_{N}\big{)}\bigg{(}(I_{N}\otimes S)\xi-\sum_{r\in\mathscr{L}}(\Phi_{r}\otimes I_{N})\hfill\\ \times\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi-\sum_{r\in\mathscr{L}}(\Phi_{r}\otimes I_{N})\gamma^{ol}\bigg{)}+\frac{1}{2}\xi^{\top}\hfill\\ \times\big{(}\operatorname{diag}(\dot{\vartheta}_{i})\otimes I_{N}\big{)}\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes I_{N}\big{)}\xi\hfill\\ \leqslant\sigma_{\max}(S)\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|\|\xi\|-\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\hfill\\ \times\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|^{2}+\sigma_{\max}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\hfill\\ \times\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|\|\gamma^{ol}\|+\frac{1}{2}\max_{i}(\dot{\vartheta}_{i})\hfill\\ \times\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|\|\xi\|\hfill\\ =-\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|\hfill\\ \times\bigg{(}\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|-\sigma_{\max}(S)\hfill\\ /\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\|\xi\|-\sigma_{\max}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}/\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\hfill\\ \times\|\gamma^{ol}\|-\frac{1}{2}\max_{i}(\dot{\vartheta}_{i})/\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\|\xi\|\bigg{)}.\hfill\\ \end{gathered}$
Denote $\phi_{a}=\sigma_{\max}(S)/\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}$ and $\phi_{b}=\sigma_{\max}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}/\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}$, which are both positive constants. For $\dot{V}^{{}^{\prime}}\leqslant 0$, we need
$\begin{gathered}\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|-\phi_{a}\|\xi\|-\phi_{b}\|\gamma^{ol}\|\hfill\\ -\frac{1}{2}\max_{i}(\dot{\vartheta}_{i})/\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\|\xi\|\geqslant 0.\hfill\end{gathered}$ (29)
A sufficient condition to guarantee (29) is
$\begin{gathered}\big{(}\exp(\vartheta_{i})-\phi_{a}-\frac{1}{2}\max_{i}(\dot{\vartheta}_{i})/\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\big{)}\|\xi_{i}\|\hfill\\ \geqslant\phi_{b}\|\gamma^{ol}_{i}\|.\hfill\end{gathered}$ (30)
A sufficient condition to guarantee (30) is $\|\xi_{i}\|\geqslant\phi_{b}$ and $\exp(\vartheta_{i})-\phi_{a}-1/2\max_{i}(\dot{\vartheta}_{i})/\sigma_{\min}(\sum_{r\in\mathscr{L}}\Phi_{r})\geqslant\|\gamma^{ol}_{i}\|$. From Assumption 4, $\|\gamma^{ol}_{i}(t)\|\leqslant\exp(\kappa^{ol}_{i}t)$, to prove that $\exp(\vartheta_{i})-\phi_{a}-1/2\max_{i}(\dot{\vartheta}_{i})/\sigma_{\min}(\sum_{r\in\mathscr{L}}\Phi_{r})\geqslant\|\gamma^{ol}_{i}\|$, we need to prove that $\exp(\vartheta_{i})-\phi_{a}-1/2\max_{i}(\dot{\vartheta}_{i})/\sigma_{\min}(\sum_{r\in\mathscr{L}}\Phi_{r})\geqslant\exp(\kappa^{ol}_{i}t)$. Based on (10), when $\|\xi_{i}\|>\max\{\sqrt{\kappa^{ol}_{i}/q_{i}},\phi_{b}\}$, which guarantees the exponential growth of $\exp(\vartheta_{i})$ dominates all other terms, $\exists t_{1}$, such that $\forall t>t_{1}$, $\exp(\vartheta_{i})-\phi_{a}-1/2\max_{i}(\dot{\vartheta}_{i})/\sigma_{\min}(\sum_{r\in\mathscr{L}}\Phi_{r})\geqslant\exp(\kappa^{ol}_{i}t)$. Hence, we obtain $\forall t>t_{1}$,
$\dot{V}^{{}^{\prime}}\leqslant 0,\>\forall\|\xi_{i}\|>\max\{\sqrt{\kappa^{ol}_{i}/q_{i}},\phi_{b}\}.$ (31)
By LaSalle’s invariance principle [24], $\xi_{i}$ is UUB. Therefore, $\Delta_{o}$ is UUB. Next, we prove that follower-observer tracking error $\varepsilon_{i}$ is UUB. From (1), (4), (9), (17) and (14), we obtain the time derivative of (16) as
$\begin{gathered}{{\dot{\varepsilon}}_{i}}={{\dot{x}}_{i}}-{\dot{\zeta}}_{i}\hfill\\ ={A_{i}}{x_{i}}+{B_{i}}{K_{i}}{x_{i}}+{B_{i}}{H_{i}}{\zeta_{i}}-{B_{i}}{\hat{\gamma}^{a}_{i}}\hfill\\ +{B_{i}}{\gamma^{a}_{i}}-\left(A_{i}+B_{i}\Pi_{i}\right)\zeta_{i}-\operatorname{exp}(\vartheta_{i})\zeta_{i}+\Delta u_{i}-\gamma_{i}^{ol}\hfill\\ =\left(A_{i}+B_{i}K_{i}\right)\varepsilon_{i}-B_{i}\hat{\gamma}_{i}^{a}-B_{i}\gamma_{i}^{a}-\operatorname{exp}(\vartheta_{i})\xi_{i}-\gamma_{i}^{ol}+\Delta u_{i}.\hfill\end{gathered}$ (32)
From the above proof, we confirmed ${\xi_{i}}$ is UUB. Considering Assumption 2, (24) and (25), we obtain that $\beta_{i}=\operatorname{exp}(\vartheta_{i})\xi_{i}-\gamma_{i}^{ol}$ is bounded. Let ${{\bar{A}}_{i}}={A_{i}}+{B_{i}}{K_{i}}$ and ${\bar{Q}_{i}}={Q_{i}}+K_{i}^{\top}{U_{i}}{K_{i}}$. Note that ${\bar{Q}_{i}}$ is positive-definite. From (15), $P_{i}$ is symmetric positive-definite. Consider the following Lyapunov function candidate
$V_{i}=\varepsilon_{i}^{T}P_{i}\varepsilon_{i},$ (33)
and its time derivative is given by
$\begin{gathered}\dot{V}_{i}=2\varepsilon^{T}_{i}P_{i}\dot{\varepsilon}_{i}\hfill\\ =2\varepsilon_{i}^{T}P_{i}\left(\bar{A}_{i}\varepsilon_{i}+B_{i}\gamma^{a}_{i}-B_{i}\hat{\gamma}^{a}_{i}-\beta_{i}+\Delta u_{i}\right)\hfill\\ \leqslant-\sigma_{\min}\left(\bar{Q}_{i}\right)\left\|\varepsilon_{i}\right\|^{2}+2\left(\varepsilon_{i}^{T}P_{i}B_{i}\gamma^{a}_{i}-\varepsilon_{i}^{T}P_{i}B_{i}\hat{\gamma}^{a}_{i}\right)\hfill\\ -2\varepsilon_{i}^{T}P_{i}\beta_{i}+2\varepsilon_{i}^{T}P_{i}\Delta u_{i}\hfill\\ \leqslant-\sigma_{\min}\left(\bar{Q}_{i}\right)\left\|\varepsilon_{i}\right\|^{2}+2\left(\varepsilon_{i}^{T}P_{i}B_{i}\gamma^{a}_{i}-\varepsilon_{i}^{T}P_{i}B_{i}\hat{\gamma}^{a}_{i}\right)\hfill\\ +2\sigma_{\max}\left(P_{i}\right)\left\|\varepsilon_{i}\right\|\left\|\beta_{i}\right\|+2\sigma_{\max}\left(P_{i}\right)\left\|\varepsilon_{i}\right\|\left\|\Delta u_{i}\right\|.\hfill\\ \end{gathered}$ (34)
Using (18) to obtain
$\begin{gathered}\varepsilon_{i}^{\top}P_{i}B_{i}\gamma^{a}_{i}-\varepsilon_{i}^{\top}P_{i}B_{i}\hat{\gamma}^{a}_{i}\hfill\\ =\varepsilon_{i}^{\top}P_{i}B_{i}\gamma^{a}_{i}-\frac{\left\|\varepsilon_{i}^{\top}P_{i}B_{i}\right\|^{2}}{\left\|\varepsilon_{i}^{\top}P_{i}B_{i}\right\|+\exp\left(-c_{i}t^{2}\right)}\exp\left(\hat{\rho}_{i}\right)\hfill\\ \leqslant\left\|\varepsilon_{i}^{\top}P_{i}B_{i}\right\|\left\|\gamma^{a}_{i}\right\|-\frac{\left\|\varepsilon_{i}^{\top}P_{i}B_{i}\right\|^{2}}{\left\|\varepsilon_{i}^{\top}P_{i}B_{i}\right\|+\exp\left(-c_{i}t^{2}\right)}\exp\left(\hat{\rho}_{i}\right)\hfill\\ =\left\|\varepsilon_{i}^{\top}P_{i}B_{i}\right\|\big{(}\left\|\varepsilon_{i}^{\top}P_{i}B_{i}\right\|\left\|\gamma^{a}_{i}\right\|+\exp(-c_{i}t^{2})\left\|\gamma^{a}_{i}\right\|\hfill\\ -\left\|\varepsilon_{i}^{\top}P_{i}B_{i}\right\|\exp\left(\hat{\rho}_{i}\right)\big{)}/\big{(}\left\|\varepsilon_{i}^{\top}P_{i}B_{i}\right\|+\exp(-c_{i}t^{2})\big{)}.\hfill\end{gathered}$ (35)
To prove that $\varepsilon_{i}^{\top}P_{i}B_{i}\gamma^{a}_{i}-\varepsilon_{i}^{\top}P_{i}B_{i}\hat{\gamma}^{a}_{i}\leqslant 0$, we need to prove that $\left\|\varepsilon_{i}^{\top}P_{i}B_{i}\right\|\left\|\gamma^{a}_{i}\right\|+\exp(-c_{i}t^{2})\left\|\gamma^{a}_{i}\right\|-\left\|\varepsilon_{i}^{\top}P_{i}B_{i}\right\|\exp\left(\hat{\rho}_{i}\right)\leqslant 0$. Define $\upsilon_{i}=\kappa^{a}_{i}/\sigma_{\operatorname{min}}(P_{i}B_{i})$, $\omega_{i}=2\sigma_{\max}\left(P_{i}\right)\big{(}\left\|\beta_{i}\right\|+\|\Delta u_{i}\|\big{)}/\sigma_{\min}\left(\bar{Q}_{i}\right)$. Then, define the compact sets $\Upsilon_{i}\,\triangleq\,\{\|\varepsilon_{i}\|\leqslant\upsilon_{i}\}$ and $\Omega_{i}\,\triangleq\,\{\|\varepsilon_{i}\|\leqslant\omega_{i}\}$. Considering Assumption 4, we obtain that $\exp(-c_{i}t^{2})\left\|\gamma^{a}_{i}\right\|\rightarrow 0$. Hence, outside the compact set $\Upsilon_{i}=\ \{\|\varepsilon_{i}\|\leqslant\upsilon_{i}\}$, $\exists t_{1}$, such that $\varepsilon_{i}^{\top}P_{i}B_{i}\gamma^{a}_{i}-\varepsilon_{i}^{\top}P_{i}B_{i}\hat{\gamma}^{a}_{i}\leqslant 0$, $\forall t\geqslant t_{1}$; outside the compact set $\Omega_{i}$, $-\sigma_{\min}\left(\bar{Q}_{i}\right)\left\|\varepsilon_{i}\right\|^{2}+2\sigma_{\max}\left(P_{i}\right)\left\|\varepsilon_{i}\right\|\left\|\beta_{i}\right\|\leqslant 0$. Therefore, combining (34), (35) and (32), we obtain, outside the compact set $\Upsilon_{i}\cup\Omega_{i}$, $\forall t\geqslant t_{1}$,
$\dot{V}_{i}\leqslant 0.$ (36)
[…]
IV Numerical Simulations
We validate the proposed SAAR control strategies through numerical simulations of a heterogeneous autonomous MAS under EU-FDI attacks on both CIL and OL. Fig. 3 illustrates the communication topology among the agents. As seen, the autonomous MAS consists of four followers (agents 1 to 4) and four leaders (agents 5 to 8), with the following dynamics:
[…]
Figure 3: Communication topology.
We consider the following EU-FDI attack signals on the CIL and OL:
$\begin{array}[]{cc}\gamma^{a}_{1}=\begin{bmatrix}2.5e^{0.07t}\\ 1.5e^{0.04t}\\ -6.6e^{0.08t}\end{bmatrix},\gamma^{ol}_{1}=\begin{bmatrix}-1.2e^{0.10t}\\ 1.5e^{0.17t}\\ 2.7e^{0.15t}\end{bmatrix},\par\gamma^{a}_{2}=\begin{bmatrix}2.3e^{0.05t}\\ -4.7e^{0.05t}\\ 11.5e^{0.04t}\end{bmatrix},\gamma^{ol}_{2}=\begin{bmatrix}3.3e^{0.06t}\\ -2.2e^{0.15t}\\ -1.7e^{0.12t}\end{bmatrix},\\[15.00002pt] \gamma^{a}_{3}=\begin{bmatrix}3.6e^{0.10t}\\ -4.7e^{0.09t}\\ -10.2e^{0.06t}\end{bmatrix},\gamma^{ol}_{3}=\begin{bmatrix}2.8e^{0.14t}\\ -5.0e^{0.04t}\\ -1.8e^{0.08t}\end{bmatrix},\par\gamma^{a}_{4}=\begin{bmatrix}-2.9e^{0.09t}\\ 5.2e^{0.06t}\\ -7.7e^{0.07t}\end{bmatrix},\gamma^{ol}_{4}=\begin{bmatrix}-5.2e^{0.04t}\\ 2.4e^{0.13t}\\ -2.1e^{0.14t}\end{bmatrix}.\end{array}$
Select $U_{1,2,3,4}=I_{2}$, and $Q_{1,2,3,4}=3I_{2}$. The controller gain matrices $K_{i}$ and $H_{i}$ are found by solving (13)-(15).
We choose $d_{s}=0.3$ for collision avoidance. The performance of the proposed SAAR control strategies, considering these safety constraints, is demonstrated through a set of comparative simulation results contrasted with non-SAAR standard control protocols [19]. The three sub-figures in Fig. 4 illustrate the initial 3D positions of the agents and compare the system responses under standard control and SAAR control following the attacks. Both EU-FDI attacks on the OL and CIL are initiated at $t=3\,\operatorname{s}$. Fig. 4(a) depicts the initial positions of the agents, while Fig. 4(b) presents the agents’ 3D positions at $t=15.08\,\operatorname{s}$ under the standard control approach, where the containment objective is not achieved after attack initiation. In contrast, Fig. 4(c) displays the agents’ 3D positions at $t=15.08\,\operatorname{s}$ under the SAAR control approach, demonstrating that the UUB containment objective is preserved despite the attacks.
Figure 4: Leader-follower 3D motion evolution: (a) $0\,\operatorname{s}$ snapshot. (b) $15.08\,\operatorname{s}$ snapshot (after attack initiation) using the standard control. (c) $15.08\,\operatorname{s}$ snapshot (after attack initiation) using the SAAR control.
Figure 5: Containment errors: (a) using conventional control (b) using SAAR control.
Fig. 5(a) illustrates the evolution of the containment error $e_{c}$ under conventional control. The dimensional components of $e_{c}$ diverge, indicating a failure in achieving containment. In contrast, Fig. 5(b) depicts the system response under SAAR control under attack injection. The SAAR controller quickly compensates after the attacks are initiated, causing the errors to stabilize and enter a steady state. This behavior demonstrates the resilience of the SAAR controller in maintaining UUB containment performance by effectively mitigating the impact of EU-FDI attacks on both CIL and OL.
Figure 6: Pairwise distance.
Fig. 6 depicts the evolution of the distances between followers, denoted as $d_{12},d_{13},d_{14},d_{23},d_{24},$ and $d_{34}$, over time, which shows the show safe distances among followers are maintained. Fig. 5(b) and Fig. 6 show that the system successfully maintains safety and achieves the UUB containment under EU-FDI attacks on both OL and CIL, validating the effectiveness of the proposed SAAR defense strategies.
V Conclusion
This letter has formulated and solved the SAAR control problem for heterogeneous MASs to mitigate EU-FDI attacks on both OL and CIL while ensuring collision-free behaviors. An attack-resilient OL and a novel compensational signal on the CIL have been designed to address the EU-FDI attacks. A SAAR controller has then been developed by formulating and solving a QP problem integrating CBF certified collision-free safety constraints. Rigorous Lyapunov stability analysis and 3D simulations have demonstrated the SAAR controller’s effectiveness in maintaining both safety and resilience in practical autonomous MAS environments, where leader agents span safe regions to guide heterogeneous follower agents while ensuring collision avoidance.